OpenStudy (anonymous):

6x-9y=18 what are intercepts x and y

OpenStudy (anonymous):

Here

OpenStudy (er.mohd.amir):

6x-9y=18 (6x/18) - (9y/18)=18/18 x/3-y/2=1 x/3+y/(-2)=1 then x-intercept =3 y-intercept= 2 because x/a+y/b=1 then x-intercept=a y -intercept=b